40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 170 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  8303   Mon Mar 18 12:02:12 2013 AnnalisaConfigurationABSLABSL setup for g-factor measurement of PRC
Motivations
The ABSL technique has been already used in the past to measure the absolute length of the interferometer's optical cavities by means of an auxiliary laser source, as described in LIGO-P1200048-v3 and in Alberto Stochino thesis work.
Using the same technique it is possible to measure the g-factor of the power recycling cavity by measuring the cavity Transverse Mode Spacing.
 
Plan for experimental setup
The auxiliary laser is set on the POY table and is injected through the ITMY window in way to follow the same path of the POY beam. It hits the AR wedge of ITMY and is reflected back to the BS and the PRM.
 
Since the main beam is P-polarized, all the optics in the central IFO are P-polarization dependent, so it is useful to P-polarize the auxiliary beam before it enters the IFO.  
I made a mode matching calculation with a la mode script, in order to mode match the auxiliary beam waist to the waist of the main laser.
However, before ordering and installing steering optics and mode maching lenses, I'm waiting to know whether someone has an NPRO laser to install on the END table in place of the broken one, otherwise the one I'm using could be taken.
In this case a possibility could be to take the auxiliary beam from the end table with an optical fiber, but it means to use the auxiliary laser alternately to lock the arm or make a measurement of TMS. If so, a new calculation for the mode matching needs to be done.
Anyway, I hope that another laser will be found!
 
In order to phase lock the auxiliary beam with the main beam, the latter will be taken from the PSL table after the PMC through a single mode fiber, which will be brought up to the POY table. This solution results to be more reliable then taking the POY beam to phase lock the two laser, because POY is related to the locking. 
 
The signal with the beat note between the two lasers can be detected by the transmission from PR2 (POP). 
 
 
 
  8302   Fri Mar 15 16:46:52 2013 JenneBureaucracyAuxiliary lockingYend table upgrade - fast track?

In light of the Yend auxiliary laser's ill health, I think we should reconsider the possibility of changing out the Yend laser table next week.

My thinking here is that if whatever the new mode matching solution is for a replacement laser (Tara has borrowed our spare NPRO that used to sit on top of the fridge, or we could take Annalisa's) requires a rework of the table layout, we might as well put the new layout onto the new table.  So, we need to figure out what laser we will put in as the new Ygreen, and what it's waist looks like.  If it just requires a small movement of existing lenses or new lenses in similar positions to the current ones, we can keep living with our current table.  But, if the mode matching solution requires enough changes to distances / lens placement / whatever, we should think seriously about putting in the new table next week.

Here's what I would like to see happen on / before Monday:

Annalisa - Mode matching solution for new laser.  If we get the laser back from Tara, this will involve first measuring the waist, otherwise we already know the waist of the ABSL laser that Annalisa is currently using.

Annalisa and Steve - Find optics for new mode matching in the lab, or order them by Monday afternoon.

Manasa - List of every screw, washer, optic, mount, etc. that will go on the new Y end table, with a notation as to whether or not we have it in-hand, and if not, what needs to happen before we do.  Also, for things that we don't have, I'd like to see a summary of temporary solutions (e.g. keep using current mount for doubling crystal while new one is being machined).

Manasa / Annalisa / Koji - will the new mode matching solution fit within the existing layout, or do we need to redo the table layout?

  8301   Fri Mar 15 15:26:13 2013 KojiUpdateGreen LockingYarm ALS laser is funny / dying

I took a look at the laser. It is probably the LD TEC (DTEC) failure.
As the temperature of the LD (DTMP) gradually deviated from 25degCish,
the DTEC voltage also went up from 2Vish to 2.1, 2.2... 

When DTEC reaches 3V, it stopped lasing. This cools the diode a bit, and
it start lasing but repeat the above process.

I am not sure which of the head and controller has the issue.

The situation did not improve much by reducing the pumping current (ADJ: -15).

BTW, Turning on/off the noise eater did not change the situation.

I think the head/controller set should be sent out to JDSU and find how they will say.

  8300   Fri Mar 15 02:19:01 2013 JenneUpdateGreen LockingYarm ALS laser is funny / dying

I turned the laser back on around 1am.  This is still happening, although right now it is turning off more often than before, maybe every 15 seconds or so.  I am going to turn off the laser for the night.

The measured laser temperature is about 45C (I have a 25,000 count offset in the Y ALS Slow control right now....higher offset, lower temp), although the measured laser temp drops to ~43.5C when the power goes down.

  8299   Fri Mar 15 02:14:27 2013 JenneUpdateCDSSimulink linking to wrong library part

Jamie and I discovered a problem with Matlab/Simulink earlier today. 

In the end suspension models, there is a subblock (with top_names) for ALS stuff.  Inside there, we use a library part called "ALS_END".  When the model was created, it included the part ...../userapps/release/isc/c1/models/ALS_END.mdl .  However, if you open up the c1scy diagram and look in the ALS block for this part, you see the part that is in ..../userapps/release/isc/common/models/ALS_END.mdl .  Note the difference - the one we want is in the c1 directory, while the one that was created (by Jamie) for the LHO One Arm Test is in the common directory. 

If you compile the c1scy model, the RCG is using the correct library part, so the information regarding which part we want is still in there. 

However, if you delete the ALS_END part from the model, put the correct one in, save, close, then reopen the model, it once again displays the wrong model.  The right click "go to library part" option brings you to the library part that is displayed, which is currently the wrong one.  THIS IS BAD, since we could start modifying the wrong things.  You do get a warning by Matlab about the file being "shadowed", so we should take heed when we see that warning, and make sure we are getting the file we want.

We are currently running Matlab version 7.11.0.584, which is r2010b.  Step 1 will be to update Matlab to the latest version, in hopes that this fixes things.  We also should change the name of our c1 part, so that it does not have the same name as the one for the sites.  This is not a great solution since we can't guarantee that we will never choose the same names as other sites, but it will at least fix this one case.  Again, if you see the warning about "shadowed" filenames, pay attention.

  8298   Fri Mar 15 01:57:02 2013 JenneUpdateGreen LockingX arm green locked in TEM00

This work earlier today had required moving the harmonic separator back closer to its original position, so that the green could get through without clipping.  I locked the Xarm (overriding the trigger) and realigned TRX to the PD and camera.

  8297   Thu Mar 14 20:22:33 2013 JenneUpdateGreen LockingXbeat attempt

I aligned the Xgreen and PSL green to overlap on the X beat PD, and reconnected the splitter which combines the X and Y beat signals and sends them to the control room.

I've been stepping the Xend laser temperature offset in steps of 20 counts, making sure the cavity unlocks and relocks on TEM00.  So far I have not seen any beat signals for the Xarm.  I've gone from 0 to 840.

I'll be back in a few hours to keep trying, although interested parties are invited to give it a whirl. 

 

  8296   Thu Mar 14 17:31:57 2013 ManasaUpdateGreen LockingX arm green locked in TEM00

Quote:

Manasa has done some work to get the Xgreen aligned, so I'll switch to trying to find that beatnote for now.

[Jenne, Manasa]

Aligned X-arm green in TEM00.

It was difficult to get the X-arm to lock in TEM00 earlier. Even when it locked, it was just a TEM00 flash. The green was mainly bad in pitch.

I started aligning with the arms flashing in IR and it was still not possible to lock. The second trial was with the arms locked in IR; I lost the green lock when the arms were aligned for IR. I aligned by overlapping the ingoing green with the reflected green visible on the steering mirror. This got the green to lock in TEM00; but still it would only stay that way for 30 sec. Jenne pointed me to Yuta's suggestion of increasing the green refl PD gain. Once I increased the PD gain, the mode stabilized.

I went ahead and centered ALS TRX on the PD and GTRX camera (on the PSL table). ALS_TRX reads 2000+ counts.

EDIT by Jenne:  This corresponds to a power of 550uW on the PSL table, measured before the first out-of-vac steering mirror.

  8295   Thu Mar 14 16:56:16 2013 ChloeUpdate QPD Circuit Design

I have sketched out the circuit design for the QPD. However, it seems like even when using a different opamp configuration, which I talked to Eric about on Friday, space will be a problem. It may be possible to squeeze everything onto a single circuit board to fit in the QPD box but what I think is more likely is that I will need to have 2 separate circuit boards both mounted within the box, one which integrates the signal from the QPD and the other which adds/subtracts (this involves many resistors which will take up a lot of space). I will continue to think about the best design for this.

I will try to have the circuit built in the next week or so, which may be difficult since I just started finals which will take most of my time. I spent most of this week writing up an ECDL proposal for a SURF with Tara. I'll make up for whatever work I miss since I'll be here for my spring break and doing little besides working in lab.

  8294   Thu Mar 14 16:41:48 2013 JenneUpdateGreen LockingYarm ALS laser is funny / dying

Quote (elog 7002, 23July2012):

Jamie and I were doing some locking, and we found that the Yarm green wasn't locking.  It would flash, but not really stay locked for more than a few seconds, and sometimes the green light would totally disappear.  If the end shutter is open, you can always see some green light on the arm transmission cameras.  So if the shutter is open but there is nothing on the camera, that means something is wrong.

I went down to the end, and indeed, sometimes the green light completely disappears from the end table.  At those times, the LED on the front of the laser goes off, then it comes back on, and the green light is back.  This also corresponds to the POWER display on the lcd on the laser driver going to ~0 (usually it reads ~680mW, but then it goes to ~40mW).  The laser stays off for 1-2 seconds, then comes back and stays on for 1-2 minutes, before turning off for a few seconds again.

Koji suggested turning the laser off for an hour or so to see if letting it cool down helps (I just turned it off ~10min ago), otherwise we may have to ship it somewhere for repairs :( 

 This is happening again to the Yend laser.  It's been fine for the afternoon, and I've been playing with the temperature.  First I have been making big sweeps, to figure out what offset values do to the actual temperature, and more recently was starting to do a finer sweep.  Using the 'max hold' function on the 8591, I have seen the beat appear during my big sweeps.  Currently, the laser temperature measurement is at the Yend, and the RF analyzer is here in the control room, so I don't know what temp it was at when the peaks appeared.

Anyhow, while trying to reaquire lock of the TEM00 mode after changing the temperature, I find that it is very difficult (the green seems misaligned in pitch), and every minute or so the light disappears, and I can no longer see the straight-through beam on the camera.  I went down to the end, and the same symptoms of LED on the laser head turning off, power out display goes to ~40mW, are happening.  I have turned off the laser as was the solution last time, in hopes that that will fix things.

Manasa has done some work to get the Xgreen aligned, so I'll switch to trying to find that beatnote for now.

  8293   Thu Mar 14 13:38:29 2013 JenneUpdateGreen LockingPSL green cleaned up

I ran a cable to the GTRX camera.  It is now input #2.  The videoswitch script input naming is modified to match this:  Input 2 used to be "IFOPO", and is now "GTRX".  Input 28 used to be "GRNT", and is now "GTRY".  Both green trans cameras are available from the video screen.

  8292   Thu Mar 14 11:51:14 2013 JamieUpdateGeneralBeatbox upgraded with output whitening, reinstalled

Quote:

I pulled the beatbox from the 1X2 rack so that I could try to hack in some output whitening filters.  These are shamefully absent because of my mis-manufacturing of the power on the board.

Right now we're just using the MON output.  The MON output buffer (U10) is the only chip in the output section that's stuffed:

2013-03-06-195232_1060x927_scrot.png

The power problem is that all the AD829s were drawn with their power lines reversed.  We fixed this by flipping the +15 and -15 power planes and not stuffing the differential output drivers (AD8672).

It's possible to hack in some resistors/capacitors around U10 to get us some filtering there.  It's also possible to just stuff U9, which is where the whitening is supposed to be, then just jump it's output over to the MON output jack.  That might be the cleanest solution, with the least amount of hacking on the board.

I modified the beatbox according to this plan.  I stuffed the whitening filter stage (U9) as indicated in the schematic (I left out the C26 compensation cap which, according to the AD829 datasheet, is not actually needed for our application).  I also didn't have any 301 ohm resistors so I stuffed R18 with 332 ohm, which I think should be fine.

Instead of messing with the working monitor output that we have in place, I stuffed the J5 SMA connector and wired U9 output to it in a single-ended fashion (ie. I grounded the shield pins of J5 to the board since we're not driving it differentially).  I then connected J5 to the I/Q MON outputs on the front panel.  If there's a problem we can just rewire those back to the J4 MON outputs and recover exactly where we were last week.

It all checks out: 0 dB of gain at DC, 1 Hz zero, 10 Hz pole, with 20 dB of gain at high frequencies.

I installed it back in the rack, and reconnected X/Y ARM ALS beatnote inputs and the delay lines.  The I/Q outputs are now connected directly to the DAQ without going through any SR560s (so we recover four SR560s). 

  8291   Thu Mar 14 04:20:54 2013 JenneUpdateGreen LockingYbeat attempt

I dedicated my evening to trying to get the Ygreen beatnote (the idea being to then get the Xgreen beatnote).

First up was tweaking up the green alignment.  Per Yuta's suggestion, elog 8283, I increased the refl PD gain by 2 clicks (20dB) to keep the lock super stable while improving the alignment.  After I finished, I turned it back to its nominal value.  I discovered that I need lenses in front of the DC PD (for Ygreen, and I'm sure Xgreen will be the same).  The beam is just barely taking up the whole 2mm diode, so beam jitter translates directly to DC power change measured by the diode.  I ended up going just by the green transmission camera for the night, and achieved 225uW of Ygreen on the PSL table.  This was ~2,000 counts, but some of the beam is always falling off the diode, so my actual counts value should be higher after installing a lens. 

I then opened up the PSL green shutter, which is controlled by the button labeled "SPS" on the shutter screen - I will fix the label during some coffee break tomorrow.  Using my convenient new PSL green setup, removing the DC PD allows the beam to reflect all the way to the fuse box on the wall, so you can check beam overlap between the PSL green and the arm green at a range of distances.  I did this for Ygreen, and overlapped the Ygreen and PSL green. 

I checked the situation of the beat cabling, since Jamie has the beatbox out for whitening filter modifications tonight.  In order to get some signal into the control room, I connected the output of the BBPD amplifier (mounted on the front of the 1X2 rack) directly to the cable that goes to the control room.  (As part of my cleanup, I put all the cables back the way I found them, so that Jamie can hook everything back up like normal when he finishes the beatbox.) 

I then started watching the signal on the 8591E analyzer, but didn't magically see a peak (one can always hope....).

I decided that I should put the offset in the Y AUX laser slow servo back to the value that we had been using for a long time, ~29,000 counts.  This is where things started going south.  After letting that go for a minute or two, I thought to go check the actual temperature of the laser head.  The "T+" temperature on the controller read something like 42C, but the voltmeter which reads a voltage proportional to the temp (10C/V) was reading 5.6V.  I immediately turned off the offset, but it's going to take a while for it to cool down, so I'll come back in the morning.  I want the AUX laser to be something like 34C, so I just have to wait.  Ooops.

Still to do (for the short-term FPMI):

* Find Y beatnote.

* Align Xgreen to the arm - it's still off in pitch.

* Align Xgreen and PSL green to be overlapped, hitting the BBPD.

* Find the X beatnote.

* Reinstall the beatbox.

* Use ALS to stabilize both arms' lengths.

* Lock MICH with AS.

* Look at the noise spectrum of AS - is there more noise than we expect (Yuta and Koji saw extra noise last summer), and if so, where does it come from?  Yuta calculated (elog 6931) that the noise is much more than expected from just residual arm motion.

* Write a talk.

  8290   Wed Mar 13 21:04:37 2013 JenneUpdateGreen LockingPSL green cleaned up

Both X and Y green are aligned such that the arm beams hit the broadband PD.  Also, the 4th port of the combining BS for each arm was used to put a camera and DC PD for each arm.  So, ALS-TRX and ALS-TRY are both active right now.  The camera currently labeled "GRNT" is the Ygreen transmission.  I have a camera installed for Xgreen transmission, but I have not run a cable to the video matrix.  For now, to speed things up, I'll just use the GRNT cable and move it back and forth between the cameras.

  8289   Wed Mar 13 16:29:25 2013 ManasaUpdateLockingTWO arms TWO colors

Quote:

- If X green lock is not tight, maybe temporarily increasing loop gain helps. This can be done by increasing the amplitude of the frequency modulation or increasing green refl PD gain. Also, if X green beam spot is too wiggly compared with Y green, it is maybe because of air flow from the air conditioner (elog #6849). I temporarily turned it off when I did X green steering last summer.

- Did you have to re-align TRX path? We moved the harmonic separator on X end table horizontally to avoid IR TRX clipping before beam centering on X arm (elog #8162). I wonder what is the current situation after the beam centering.

 - We tried locking with the air conditioning switched off at the X-end

- TRX path is unaltered (IR still goes through the center of the harmonic separator.

  8288   Wed Mar 13 16:25:16 2013 ManasaUpdateIOOTuning MC length/FSR

[Koji, Manasa]

We looked at the MC modulation frequency on the spectrum analyzer and observed beat notes between MC modulation freq (29.5MHz) and modulation frequencies (11.06 MHz and 55.3MHz).

Beat notes were suppressed by changing the carrier frequency from 11.065910 MHz to 11.066147MHz. 

Detailed discussion and data will be posted in the next elog.

  8287   Wed Mar 13 16:04:24 2013 steveUpdateSAFETYlaser safety glasses checked

 All safety glasses were cleaned in soapy water by Bob. I measured their transmission at 1064 nm, 150 mW,  beam diameter 1.5 mm  They are in working order, no transmission.

 Inventory:

10 pieces of KG-5, fit over, from Laser Safety

 4 pieces of KG-5, std size, from Drever Lab, best visibility

 1 piece of KG-5 coated for visible, std size, from Kentek

15 pieces of green-plastic LOTG-YAG, fit over, from UVEX

 7 pieces of green-plastic B-D+S 137, std areo fit, from Sperian

 3 pieces of green-plastic, old Thorlab, fit over

 2 pieces of green-plastic, fit over, from Laservision

 8 pieces of braun- plastic, fit over, for green & IR protection, from UVEX & Thorlabs

Attachment 1: safetyglasseschecked.jpg
safetyglasseschecked.jpg
  8286   Wed Mar 13 15:30:37 2013 Max HortonUpdateSummary PagesFixing Plot Limits

Jamie has informed me of numpy's numpy.savetxt() method, which is exactly what I want for this situation (human-readable text storage of an array).  So, I will now be using:

      # outfile is the name of the .png graph.  data is the array with our desired data.
      numpy.savetxt(outfile + '.dat', data)

to save the data.  I can later retrieve it with numpy.loadtxt()

  8285   Wed Mar 13 11:34:24 2013 SteveUpdateoptical tablestable covers moved to CES

The two acrylic optical table enclosures were moved from the carpenter shop to CES. I need to order windows. The latest quotes from Laseroptik  are posted at the wiki / aux_optics page.

Things to do: order windows, draw and order window flange, install surgical tubing seals, buy and line enclosures with IR shield films.

Attachment 1: acrylicEnclosure.jpg
acrylicEnclosure.jpg
  8284   Wed Mar 13 10:26:58 2013 ManasaUpdateLockingWe're still good!!

We're still good with the IFO alignment after 7hours.

I found the green still locked in the same state as last night; but no IR (so the arms are stable and the TTs should definitely take the blame).

From last night's observation (elog about drift in TT1), I only moved TT1 in pitch and gained back locking in IR for both the arms

  8283   Wed Mar 13 08:34:33 2013 yutaUpdateLockingTWO arms TWO colors

- I took the shutter from AS table to use it for the PSL green. It was sitting near MC REFL path unused (elog #8259).

- If X green lock is not tight, maybe temporarily increasing loop gain helps. This can be done by increasing the amplitude of the frequency modulation or increasing green refl PD gain. Also, if X green beam spot is too wiggly compared with Y green, it is maybe because of air flow from the air conditioner (elog #6849). I temporarily turned it off when I did X green steering last summer.

- X green transmission on PSL table reached ~270 uW last summer (elog #6849, elog #6914). Y green transmission is now ~240 uW and ~2700 counts at maximum. So, X green transmission should reach ~3000 in counts.

- Did you have to re-align TRX path? We moved the harmonic separator on X end table horizontally to avoid IR TRX clipping before beam centering on X arm (elog #8162). I wonder what is the current situation after the beam centering.

  8282   Wed Mar 13 03:12:47 2013 Manasa, JenneUpdateLockingTWO arms TWO colors

[Jenne, Manasa]

2 colors 2 arms realized!

1. Spot centering:

We spot centered the IR in both arms.
- Use TT1 and TT2 to center in Y arm (I visually center the spots on the ITM and ETM and then use TTs iteratively)
- Use BS-ETM to center in X arm

Spot positions after centering
               X arm            Y arm
         itmx    etmx        itmy    etmy
pitch    -0.86    0.37        1.51    0.05
yaw      0.01    -0.1        0.08    0.10


2. TT1 drifting in pitch (Bistable)
During the arm alignment routine for spot centering, we observed that TRY dropped (from TRY = 0.9 until the arm lost lock) every 40minutes or so. The arm was relocked by moving TT1 in pitch. The (locking - unlocking due to drift - relocking) cycle was monitored and we observed that it was bistable i.e. if TT1 was moved up in pitch (0.2 on the slider) to relock for the first time ; the next time it lost lock, TT1 had to be moved down by nearly the same distance to relock the arm.
Moving TT2 or the testmasses did not help with relocking the arms; so TT1 seems to be the one causing all th trouble atleast for today.

3. ALS - green alignment

We then moved on to Ygreen.  We used the out of vac steering mirrors to center the beam on the 2 irises that are in place on the table, which was a good starting place.  After doing that, and tweaking a small amount to overlap the incident and reflected beams on the green steering mirrors, we saw some mode lock.  We adjusted the end table steering mirrors until the Ygreen locked on TEM00.  We then followed Rana's suggestion of locking the IR to keep the cavity rigid while we optimized the green transmission.  Yuta, while adjusting ITMY and ETMY (rather than the out of vac mirrors) had been able to achieve a green transmission for the Yarm of ~2700 counts using the GTRX DC PD that's on the table. We were only able to get ~2200, with brief flashes up to 2500.

After that, we moved on to the X arm.  Since there are no irises on the table, we used the shutter as a reference, and the ETM optic itself.  Jenne looked through the viewport at the back of the ETM, while Manasa steered mirrors such that we were on the center of the ETM and the shutter.  After some tweaking, we saw some higher order modes lock.  We had a very hard time getting TEM00 to stay locked for more than ~1 second, even if the IR beam was locked.  It looks like we need to translate the beam up in pitch.  The leakage of the locked cavity mode is not overlapped with the incident beam or the promptly reflected beam.  This indicates that we're pretty far from optimally aligned.  Manasa was able to get up to ~2000 counts using the same GTRX PD though (with the Ygreen shutter closed, to avoid confusion).  Tomorrow we will get the Xarm resonating green in the 00 mode.

We need to do a little cleanup on the PSL green setup.  Yuta installed a shutter (I forget which unused one he took, but it was already connected to the computers.), so we can use it to block the PSL green beam.  The idea here is to use the 4th port of the combining beam splitters that are just before each beat PD, and place a PD and camera for each arm.  We already have 2 PDs on the table connected to channels, and one camera, so we're almost there. Jenne will work on this tommorrow during the day, so that we can try to get some beat signals and do some handoffs in the evening.

  8281   Wed Mar 13 02:48:13 2013 JenneUpdateIOOMC all better

Koji reminded me (again....this is probably the  2nd or 3rd time I've "discovered" this, at least) that the script

..../scripts/MC/WFS/WFS_FilterBank_offsets

exists, and that we should use it sometimes.  See his elog 7452 for details. 

 

Notes about using this script:

* Only use it after MC has been very well aligned.  MC REFL DC should be less than 0.5 when the MC is locked (with the DC value ~4.5 with the MC unlocked, as usual).  This is hard to achieve, but important.  Also, check the MC spot centering.

* With the WFS servo off, but the MC locked and light on the WFS diodes, run the script. 

  8280   Tue Mar 12 14:51:00 2013 SteveUpdateComputersbuy warranty or not ?

 Details of the warranties are posted on wiki power supply cost, warranty described, cost

.......I’ve also attached a warranty renewal quote.  A 1 year warranty renewal is usually $.... per year, but we gave you special pricing of $.... / year if you renew both units.  This pricing is also special due to the fact that both warranties expired awhile ago.  We usually require that the warranty renewal begin on the date of expiration, but we will waive this for you this time if both are renewed.

 

JetStor SATA 416S, SN: SB09040111A3 – expired 04/24/2012 (3 years old)

 

JetStor SATA 516F, SN: SB09080016P – expired on 08/21/2012........

 

. Are we keep it for an other 2 years? buy warranty or buy better storage.

 

  8279   Tue Mar 12 14:02:22 2013 JenneUpdateAlignmentBeam drift - mystery partially solved?

Steve just told those of us in the control room that the custodian who goes into the IFO room regularly steps on the blue support beams to reach the top of the chambers to clean them.  Since we have seen in the past that stepping on the blue tubes can give the tables a bit of a kick, this could help explain some of the drift, particularly if it was mostly coming from TT2.  The custodian has promised Steve that he won't step on the blue beams anymore.

This doesn't explain any of the ~1 hour timescale drift that we see in the afternoons/evenings, so that's still mysterious.

  8278   Tue Mar 12 12:06:22 2013 JamieUpdateComputersFB recovered, RAID power supply #1 dead

The framebuilder RAID is back online.  The disk had been mounted read-only (see below) so daqd couldn't write frames, which was in turn causing it to segfault immediately, so it was constantly restarting.

The jetstor RAID unit itself has a dead power supply.  This is not fatal, since it has three.  It has three so it can continue to function if one fails.  I have removed the bad supply and gave it to Steve so he can get a suitable replacement.

Some recovery had to be done on fb to get everything back up and running again.  I ran into issues trying to do it on the fly, so I eventually just rebooted.  It seemed to come back ok, except for something going on with daqd.  It was reporting the following error upon restart:

[Tue Mar 12 11:43:54 2013] main profiler warning: 0 empty blocks in the buffer

It was spitting out this message about once a second, until eventually the daqd died.  When it restarted it seemed to come back up fine.  I'm not exactly clear what those messages were about, but I think it has something to do with not being able to dump it's data buffers to disk.  I'm guessing that this was a residual problem from the umounted /frames, which somehow cleared on it's own.  Everything seems to be ok now.

Quote:

Manasa just went inside to recenter the AS beam on the camera after our Yarm spot centering exercises of the evening, and heard a loud beeping.  We determined that it is the RAID attached to the framebuilder, which holds all of our frame data that is beeping incessantly.  The top center power switch on the back (there are FOUR power switches, and 3 power cables, btw.  That's a lot) had a red light next to it, so I power cycled the box.  After the box came back up, it started beeping again, with the same front panel message:

H/W monitor power #1 failed.

DO NOT DO THIS.  This is what caused all the problems.  The unit has three redundant power supplies, for just this reason.  It was probably continuing to function fine.  The beeping was just to tell you that there was something that needed attention.  Rebooting the device does nothing to solve the problem.  Rebooting in an attempt to silence beeping is not a solution.  Shutting of the RAID unit is basically the equivalent of ripping out a mounted external USB drive.  You can damage the filesystem that way.  The disk was still functioning properly.  As far as I understand it the only problem was the beeping, and there were no other issues.  After you hard rebooted the device, fb lost it's mounted disk and then went into emergency mode, which was to remount the disk read-only.  It didn't understand what was going on, only that the disk seemed to disappear and the reappear.  This was then what caused the problems.  It was not the beeping, it was the restarting the RAID that was mounted on fb.

Computers are not like regular pieces of hardware.  You can't just yank the power on them.  Worse yet is yanking the power on a device that is connected to a computer.  DON"T DO THIS UNLESS YOU KNOW WHAT YOU"RE DOING.  If the device is a disk drive, then doing this is a sure-fire way to damage data on disk.

 

  8277   Tue Mar 12 11:49:18 2013 JenneUpdateIOOMC weird again

[Manasa, Annalisa, Jenne]

The MC wasn't locking on TEM00 this morning, and the WFS kept pulling the MC out of alignment.  The MC was realigned, and the WFS spots are back to being roughly centered (all of this only touching the MC sliders), but the WFS keep doing bad things.  They're okay, and improve the alignment slightly at first, but as soon as the FM1 integrator comes on, the MC alignment immediately starts going bad, and within a second or so the MC has unlocked.

The WFS are off right now, and we'll keep investigating after LIGOX.

  8276   Tue Mar 12 00:58:05 2013 ManasaUpdateAlignmentYarm - Spot positions centered

[Jenne, Manasa]


Spot centering on Y arm - DONE!

Alignment procedure
1. I went back to the IFO alignment slider positions from Friday. The Y arm was flashing in HOM because the earthquake this morning tripped all suspensions and the slider values were not real. X arm did not have any flashes.

2. Y arm aligned using TT1 and TT2. Spot centering measured using Jenne's A2L_Yarm script.

Spot positions:
           ITMY    ETMY
Pitch    6.48    4.39
Yaw     -7.42    -3.135

3. I started centering in pitch. I used the same in-vac alignment method (down on TT1 and up on TT2 in pitch) and measured spot positions.

4. When the spot positions were centered in pitch, I started with yaw alignment.

5. I had to use TT1 to center on ITMY and move TT2 and ITMY to center on ETMY.

6. Spot positions after centering:

                          ITMY    ETMY
Pitch    -1.22    -1.277
Yaw       0.42    -0.731


7. I wanted to go back and tweak the pitch cenetering; but framebuilder failed and dataviewer kept loosing connection to fb

Notes
AS seems clipped. Although it could be because of the misaligned BS.

IPANG was centered on the QPD, but it is so clipped, that I'm not sure we can trust it.  Max sum right now is -4, rather than the usual -8 or -9.

Tomorrow:

Once fb is fixed, we should align the X-arm which will be followed by green alignment.

Mystery
Over the last few weeks, it has been observed that there is some strong seismic activity that starts at around 9PM everyday and goes on for a couple of hours. It seems unlikely that it is our geologist neighbour (Jenne met with the grad student who works on the noisy experiment).
 

  8275   Tue Mar 12 00:45:50 2013 JenneUpdateIOOMC weird again

~20 minutes ago, maybe right around the time the fb's RAID died (elog 8274) the mode cleaner started behaving weirdly again.  The reflected value is very high, even with the WFS on.  Earlier this evening, I saw that with the WFS off, the MC reflection was high, but the WFS brought it back down to ~0.7 or 0.8.  But now it's ~1.3.  With the WFS off, the reflected value is ~1.1.  I don't really understand.

Also, the PMC has been drifting in alignment in pitch all day, but a lot more later in the day. The PMC trans is 0.800 right now, but it was as high as 0.825 today, and spent most of the day in the high 0.81xxx range today.

I would provide plots, but as mentioned in elog 8274, we can't get data right now.

  8274   Tue Mar 12 00:35:56 2013 JenneUpdateComputersFB's RAID is beeping

[Manasa, Jenne]

Manasa just went inside to recenter the AS beam on the camera after our Yarm spot centering exercises of the evening, and heard a loud beeping.  We determined that it is the RAID attached to the framebuilder, which holds all of our frame data that is beeping incessantly.  The top center power switch on the back (there are FOUR power switches, and 3 power cables, btw.  That's a lot) had a red light next to it, so I power cycled the box.  After the box came back up, it started beeping again, with the same front panel message:

H/W monitor power #1 failed.

Right now the fb is trying to stay connected to things, and we can kind of use dataviewer, but we lose our connection to the framebuilder every ~30 seconds or so.  This rough timing estimate comes from how often we see the fb-related lights on the frontend status screen cycle from green to white to red back to green (or, how long do the lights stay green before going white again).  We weren't having trouble before the RAID went down a few minutes ago, so I'm hopeful that once that's fixed, the fb will be fine. 

In other news, just to make Jamie's day a little bit better, Dataviewer does not open on Pianosa or Rosalba.  The window opens, but it stays a blank grey box.  This has been going on for Pianosa for a few days, but it's new (to me at least) on Rosalba.  This is different from the lack of ability to connect to the fb that Rossa and Ottavia are seeing.

  8273   Mon Mar 11 22:28:30 2013 Max HortonUpdateSummary PagesFixing Plot Limits

Quick Note on Multiprocessing:  The multiprocessing was plugged into the codebase on March 4. Since then, the various pages that appear when you click on certain tabs (such as the page found here: https://nodus.ligo.caltech.edu:30889/40m-summary/archive_daily/20130304/ifo/dc_mon/ from clicking the 'IFO' tab) don't display graphs.  But, the graphs are being generated (if you click here or here, you will find the two graphs that are supposed to be displayed).  So, for some reason, the multiprocessing is preventing these graphs from appearing, even though they are being generated.  I rolled back the multiprocessing changes temporarily, so that the newly generated pages look correct until I find the cause of this.

Fixing Plot Limits:  The plots generated by the summary_pages.py script have a few problems, one of which is: the graphs don't choose their boundaries in a very useful way.  For example, in these pressure plots, the dropout 0 values 'ruin' the graph in the sense that they cause the plot to be scaled from 0 to 760, instead of a more useful range like 740 to 760 (which would allow us to see details better).

The call to the plotting functions begins in process_data() of summary_pages.py, around line 972, with a call to plot_data().  This function takes in a data list (which represents the x-y data values, as well as a few other fields such as axes labels).  The easiest way to fix the plots would be to "cleanse" the data list before calling plot_data().  In doing so, we would remove dropout values and obtain a more meaningful plot.

To observe the data list that is passed to plot_data(), I added the following code:

      # outfile is a string that represents the name of the .png file that will be generated by the code.
      print_verbose("Saving data into a file.")
      print_verbose(outfile)
      outfile_mch = open(outfile + '.dat', 'w')

      # at this point in process_data(), data is an array that should contain the desired data values.
      if (data == []):
          print_verbose("Empty data!")
      print >> outfile_mch, data
      outfile_mch.close()

When I ran this in the code midday, it gave a human-readable array of values that appeared to match the plots of pressure (i.e. values between 740 and 760, with a few dropout 0 values).  However, when I let the code run overnight, instead of observing a nice list in 'outfile.dat', I observed:

[('Pressure', array([  1.04667840e+09,   1.04667846e+09,   1.04667852e+09, ...,
         1.04674284e+09,   1.04674290e+09,   1.04674296e+09]), masked_array(data = [ 744.11076965  744.14254761  744.14889221 ...,  742.01931356  742.05930208
  742.03433228],
             mask = False,
       fill_value = 1e+20)
)]

I.e. there was an ellipsis (...) instead of actual data, for some reason.  Python does this when printing lists in a few specific situations.  The most common of which is that the list is recursively defined.  For example:

INPUT:
a = [5]
a.append(a)
print a

OUTPUT:
[5, [...]]

It doesn't seem possible that the definitions for the data array become recursive (especially since the test worked midday).  Perhaps the list becomes too long, and python doesn't want to print it all because of some setting.

Instead, I will use cPickle to save the data.  The disadvantage is that the output is not human readable.  But cPickle is very simple to use.  I added the lines:

      import cPickle
      cPickle.dump(data, open(outfile + 'pickle.dat', 'w'))

This should save the 'data' array into a file, from which it can be later retrieved by cPickle.load().

Next:
There are other modules I can use that will produce human-readable output, but I'll stick with cPickle for now since it's well supported.  Once I verify this works, I will be able to do two things:
1) Cut out the dropout data values to make better plots.
2) When the process_data() function is run in its current form, it reprocesses all the data every time.  Instead, I will be able to draw the existing data out of the cPickle file I create.  So, I can load the existing data, and only add new values.  This will help the program run faster.

  8272   Mon Mar 11 19:01:21 2013 JenneUpdatePEMproposed seismometer locations

This is my interpretation of where Steve is proposing to place the seismometers (he wrote ITMX southwest, but I'm pretty sure from the photo he means southeast).

I think his point is that these locations are on the less-used side of the beam tube, so they will not be in the way.  Also, they are not underneath the tube, so we will not have any problems putting the covers on/taking them off.

 

 

SeismometerProposedPositions_11March2013.pdf

  8271   Mon Mar 11 17:18:00 2013 ManasaUpdateEnvironmentEarthquake: Suspensions tripped and MC realigned

I found all suspensions including the MC suspensions tripped this morning after the earthquake.

I damped all the optics and realigned MC mirrors to lock at refl 0.57.

PRM and SRM tripped a couple of times due to the aftershocks that followed; but were damped eventually.

  8270   Mon Mar 11 16:29:29 2013 SteveUpdatePEMproposed seismometer locations

Granite base 20" x 20" x 5"  locations are on the CES side of our IFO arms:  as shown ETMY_ south-west, ETMX_north-east, ITMX_south-east . No height limitation. This side of the tube has no traffic.

SS cover McMaster#  41815T4  (H) SS container cov

Attachment 1: ETMY_sw.jpg
ETMY_sw.jpg
Attachment 2: ITMX_se.jpg
ITMX_se.jpg
Attachment 3: ETMX_ne.jpg
ETMX_ne.jpg
  8269   Mon Mar 11 15:52:46 2013 JenneUpdateIOOMC spots centered, WFS realigned

Looking more into the MC, it appears that no spot centering was done after the power attenuation optics were removed (see elog 8142).  Since Manasa had changed the zig-zag steering after putting in the attenuation optics (elog 7843) the pointing was not correct for the nominal MC.  This is (likely) why Yuta and Manasa found some significant decentering.  If Manasa's tweaks when preparing for vent were primarily in yaw, then this is most definitely what happened.  A note, this is why we should change to inserting the attenuation optics before the PMC, but even so, one should adjust the angle of the PBS, NOT the zig zag optics, to get the input pointing back to the same place.  This is also why it is useful to ensure the attenuation optics do not block the PSL pointing QPD pickoff, so it is easier to adjust the PBS to get back to the original pointing.

In any case, I touched the last zig-zag mirror in yaw today (the top of the knob was moved away from me) to recenter the spots.

MCspots_11Mar2013.png

With the MC unlocked, I centered the WFS.  Now the MC is back to its normal working condition....WFS are on, autolocker is on, reflected DC power is low, etc., etc.

  8268   Mon Mar 11 14:10:05 2013 JenneUpdateEnvironmentMC suspensions moved by this morning's earthquakes

None of the suspensions All suspensions were tripped (edited by Manasa; see elog 8271) by this morning's earthquakes, but the MC suspensions are in a different place than they were a day ago.  The big symptom here is that the MCWFS are pulling the mode cleaner slightly out of alignment.  When it first locks, the reflected light is ~0.5, but when the WFS are engaged it goes up to ~0.8.  I'm going to put the MC optics back where they were (according to SUSpit and SUSyaw), and tweak up the MC from there. Probably other optics are affected, but I was going to work on continuing to center the beam on the Yarm optics, so I'll deal with the rest of the IFO in a minute.

Note re: lower plot - the mxstream was down on c1sus and c1ioo, so no fast channels on those computers were recorded for almost a day.  (The plot is one day 4 days long).  I was going to plot the seismic blrms along with the suspension pointing values, but there's no data saved, so there's no point.  Jamie tells me he thinks this spontaneous loss of the mxstream is fixed in the next RCG upgrade, and that we can talk about upgrading the RCG after the LSC meeting, so this data loss is no longer an issue.

EQ.png

MC-SUS-kicked-EQ.png

EDIT:  Plot with 4 days of trend, rather than just 1.  The MC alignment (as measured by MC refl) has been very bad for several days.  I'm going to move the suspensions back to their last good place.  Also, Manasa realigned the MC after the EQ, so I don't actually know where the suspensions got kicked to this morning.

  8267   Mon Mar 11 12:29:25 2013 ManasaUpdateAlignmentInput beam drift - weekend trend

I centered ipang and ippos on the QPDs (using only the steering mirrors) and wanted to see the drift over the weekend.

Observation
1. IPANG has drifted (QPD sum changed from -6 to -2.5); but it is still on the QPD.
2. IPPOS does not show any drift.
3. In the plot: The jump in IPANG on the left occured when I centered the beam to the QPD and that on the right is from the 4.7 earthquake and its aftershocks this morning.

Follow-up questions
1. Do we need to worry about this drift?
2. Which of the two TTs is resposible for the drift?
3. Do the TTs tend to drift in the same direction everytime?

P.S. The TTs were not touched to center on IPANG and IPPOS. The last time they were touched was nearly 6 hours before the centering. So the question of any immediate hysteresis is ruled out.

IPANG_drift.png

  8266   Mon Mar 11 10:20:36 2013 Max HortonSummaryComputersAttempted Smart UPS 2200 Battery Replacement

Attempted Battery Replacement on Backup Power Supply in the Control Room:

I tried to replace the batteries in the Smart UPS 2200 with new batteries purchased by Steve.  However, the power port wasn't compatible with the batteries.  The battery cable's plug was too tall to fit properly into the Smart UPS port.  New batteries must be acquired.  Steve has pictures of the original battery (gray) and the new battery (blue) plugs, which look quite different (even though the company said the battery would fit).

The Correct battery connector is GRAY : APC RBC55

Attachment 1: upsB.jpg
upsB.jpg
Attachment 2: upsBa.jpg
upsBa.jpg
  8265   Sun Mar 10 13:29:29 2013 KojiHowToIOOHow to calculate the accumulated round-trip Gouy phase

How to calculate the accumulated round-trip Gouy phase (and namely the transverse mode spacing) of a general cavity
only from the round-trip ABCD matrix

T1300189

  8264   Sat Mar 9 19:29:27 2013 KojiUpdatePSLModulation depth

Last night I measured the modulation depth of the MC incident beam.


Method:

The beam is taken from one of the  PO beam at the wedge plate before the IMC.
After removing the knife edge to dump this beam, the beam is sent to the west side
of the PSL table and put into the OSA cavity.
[The beam dump was returned after the measurement.]

I had some confusion and after all I use the OSA labeled as AS OSA rather than the one on the PSL table.
[The AS OSA was returned to the AP table.]

The transmission was detected by PDA255 and filtered by ITHACO 1201 preamp with G=10, no HPF, 30kHz LPF.
It was confirmed that the peak amplitudes are not reduced by the LPF filter. The resulting time series
was recorded by an oscilloscope.

Three measurements have been taken. The 11MHz peaks are offset by the carrier peak. They appropriately
removed. The ratio of the sideband and carrier peaks is converted to the modulation depth using the following formula.

P_sb / P_ca = [J1(m)/J0(m)]^2


Measurement

The modulation depth for the 11MHz: 0.190 +/- 0.003

The modulation depth for the 55MHz: 0.2564 +/- 0.0003

The three scans showed very similar numbers. That's why the statistical error is such small.
I don't think the systematic error is not such good.

This number is much different form the previous meaurement by Mirko.

http://nodus.ligo.caltech.edu:8080/40m/5519 m=0.14 (11MHz) & 0.17 (55MHz)
but the measured voltages and the modulatio depths are inconsistent.

http://nodus.ligo.caltech.edu:8080/40m/5462 m=0.17 (11MHz) & 0.19 (55MHz)

Probably the modulation depths should be checked by the IMC again.
However, it is certain that the 55MHz modulation exists, and even larger than the 11MHz one.

The next is to confirm that the modulation frequency is matched with the IMC FSR.
It is to make sure that the modulation is transmitted to the main IFO without attenuation.

Attachment 1: mod_depth.pdf
mod_depth.pdf
  8263   Fri Mar 8 21:00:05 2013 ManasaUpdatePSLPMC fixed

Quote:

  Mystery

1. Filter module (FM1) on PRCL and MICH show significant delay while enabling and disabling.

2. I tried to fix PMC alignment (PMC trans was 0.76). I was not able to get PMC trans more than 0.79.
PMC has been this way since yesterday.

3. MICH is still bright when locked (ASDC_OUT reads 0.08 for dark and 2.0 for bright). We suspect it is because of the AS55_I error offset that persists even after running LSCoffsets script.

4. PRMI shows some dither at 3Hz when locked.

 [Koji, Manasa]

PMC is fixed with 0.84 in transmission. It was misaligned in pitch (fixing this increased PMC_trans to 0.822 from 0.773) and Koji also touched the wave plate and polarizer (changed PMC_trans to 0.845).

  8262   Fri Mar 8 20:51:00 2013 ManasaUpdateAlignmentInput beam drift - investigation **IMPORTANT**

Checking the drift in input pointing (TT2 is the main suspect)

I have centered IPPOS and the 2/3 part of IPANG that comes out of vacuum to the QPDs to see the drift in input pointing over the weekend or atleast overnight.

If anybody would be working with the IFO alignment over the weekend, do so only after recording the drift in IPANG and IPPOS or if you will be working later tonight, center them ion the QPDs before leaving.

  8261   Fri Mar 8 16:05:56 2013 yutaBureaucracyGeneralaction items for PRMI / ALS-FPMI

We should focus our work both on PRMI and ALS-FPMI (elog #8250).

CDS:

    - Check out ASS and A2L working -JENNE (ALS done, ASS on going elog #8229)
    - Are all whitening filters for PDs toggling correctly? -JENNE, JAMIE (POX11 was OK, elog #8246)

PRMI locking:
    - Adjust I/Q rotation angles for error signals -JENNE, YUTA (coarsely done elog #8212)
    - Adjust filters -JENNE, YUTA (coarsely done elog #8212)
    - Coil balancing for BS (and ITMs/ETMs) -YUTA (done elog #8182)
    - Calculate sensing matrix for PRMI and convert them into physical units -JENNE, JAMIE
    - Measure sensing matrix for PRMI -JENNE, MANASA
    - Measure 55 MHz modulation depth -KOJI

PRC characterization in PRMI:

    - Measure PR2 loss from flipping -MANASA (on going elog #8063)
    - Measure mode matching ratio -JENNE, YUTA
    - Measure finesse, PR gain -JENNE, YUTA (done elog #8212)
    - Calibrate PRM and/or ITM oplevs -MANASA, YUTA (done elog #8221)
    - Measure g-factor by tilting PRM or ITMs -JAMIE, YUTA (coarsely done elog #8235, use other methods to check)
    - Simulate intra-cavity power dependance on PRM tilt -JAMIE (see elog #8235)
    - Calculate expected finesse, PR gain -JENNE
    - Mode match and align aux laser from POY -ANNALISA (on going elog #8257)

ALS:
    - Prepare for installation of new end tables on next vent -MANASA
    - Install green DC PDs and cameras on PSL table -JENNE, MANASA
    - Make ALS handing off to DARM/CARM LSC script -JENNE, YUTA
    - Demonstrate FPMI using ALS -JENNE, YUTA
    - Phase tracker characterization -YUTA, KOJI (bad whitening elog #8214)
    - better beatbox with whitening filters -JAMIE, KOJI

Others:
    - Update optical layout CAD after PR2 flipping -MANASA
    - IMC REFL demod phase rotation -EVAN, ANNALISA (done elog #8185)
    - Look into PMC drift -JENNE, MANASA
    - Measure RFAM contribution to error signals -YUTA
    - Look into TT2 drift -JENNE, MANASA

  8260   Fri Mar 8 16:02:52 2013 JenneUpdateAlignmentGetting closer to beam centering on Yarm

I'm working on getting the input beam centered on the Yarm optics.  To do this, I measured the spot positions, move the tip tilts, realign the cavity, then measure the new spot positions.  While doing this, I am also moving the BS and Xarm optics to keep the Xarm aligned, so that I don't have to do hard beam-finding later.

Here is the plot of spot measurements today.  The last measurement was taken with no moving, or realigning, just several hours later after speaking with our Indian visitors.  I'm closer than I was, but there is more work to do.

YARMdecenter_zoom_8Mar2013.png

  8259   Fri Mar 8 15:27:42 2013 yutaUpdateGreen LockingPSL green shutter installed

[Manasa, Yuta]

Mechanical shutter for PSL green is installed right in front of PSL doubling crystal.
This is for blocking PSL green when we want to measure the power of green beam from the arms.

The shutter was previously sitting on AS table un-used. Channel name to control this shutter was C1:AUX-SPS_Shutter. This should be renamed as C1:AUX-GREEN_PSL_Shutter.

Next:
  We are going to restore both arm green in parallel to PRMI work.

  - Coarsely align IR input pointing and arms using A2L
  - Align X green
  - Install green DC PDs and cameras on PSL table

  8258   Fri Mar 8 13:42:35 2013 JenneUpdateABSLBS installed on ITMY table

Re:  POY beam reduction.

We are able to lock the Yarm with the beam / gain as it is.  I had thought we might need to increase the DC gain in the whitening board by a factor of 2, but so far it's fine.

  8257   Fri Mar 8 12:57:57 2013 AnnalisaUpdateABSLBS installed on ITMY table

 Sendhil and I installed the S polarized BS on the ITMY table to steer the NPRO beam through the AR wedge and align it to the POY beam. 

We took a shutter from the BSPRM table (which was not used) and a beam dump from the AS table (which was used by the auxiliary laser already removed and installed on the ITMY).

To do: do better alignment of the NPRO beam, maybe installing some iris after the BS and before the AS wedge, phase lock the two beams. 

  8256   Fri Mar 8 03:07:19 2013 yutaUpdateLSCcalibrated PRM-ITMY length spectra

Measured free swing PRM-ITMY length was 230 nm RMS.
MI differential length was 85 nm RMS(elog #8248). This tells you that PR2, PR3 are not so noisy compared with usual suspensions.

Openloop transfer function:
  OLTF of PRM-ITMY cavity lock using REFL55_Q_ERR as error signal and PRM as actuator is below.
  UGF ~ 120 Hz, phase margin ~ 50 deg.
  Somehow, phase delay was 460 usec, which is smaller than the empirical value 550 usec.
LSCPRCLOLTF_PRITMY.png


PRM-ITMY length spectra:
  Below. Calibration was done using calibrated REFL55_Q_ERR and actuator response(elog #8255).
PRITMYmotion.png

  8255   Fri Mar 8 02:17:04 2013 yutaUpdateLSCcalibration of PRM actuator

[Manasa, Yuta]

We measured AC response of PRM actuator using PRM-ITMY cavity.
Result is

PRM:  (19.6 +/- 0.3) x 10^{-9} (Hz/f)^2 m/counts

It is almost the same as in 2011 (elog #5583). We took the same procedure as what Kiwamu did.

What we did:
  1. Aligned PRMI in usual procedure, mis-aligned ITMX and locked PRM-ITMY cavity using REFL55_Q_ERR. POP DC was about 18 when locked.

  2. Made UGF of PRM-ITMY cavity lock at 10 Hz and introduced elliptic LPF at 50 Hz(OLTF below).
OLTF_PRCL.png


  3. Measured transfer function from C1:LSC_ITMY_EXC to C1:LSC_REFL55_Q_ERR. Dividing this by ITMY actuator response(measured in elog #8242) gives calibration of REFL55_Q.

  4. Measured transfer function from C1:LSC_PRM_EXC to C1:LSC_REFL55_Q_ERR to calibrate PRM actuator.

Result:
  Calibration factor for REFL55_Q for PRM-ITMY cavity was (1.37 +/- 0.02) x 10^9 counts/m (plot below). Error is mainly from statistical error of the average.
calibREFL55Q.png


  Measured AC response (50-200 Hz) of PRM is below.
actcalibPRM.png


Next:
  - Measure free-run length spectrum of PRM-ITMY cavity and compare with MICH free-run.

  8254   Thu Mar 7 18:48:43 2013 yutaUpdateComputer Scripts / Programsreleasing my secret scripts

I released/updated my secret scripts to real scripts directory.
I checked they run correctly (but maybe not working correctly).


burtlookup.py
  in ./scripts/general/burtlookup.py

  It returns a value of a specified channel in the past using burt snapshots.
  Help is available.


GRtoggler.py
  in ./scripts/ALS/GRtoggler.py

  Toggles green shutter until it locks TEM00.
  Help is available. Threshold setting is critical.


MCbeeper.py
  in ./scripts/MC/MCbeeper.py

  Beeps when MC is unlocked.


yutalib.py
  in ./scripts/pylibs/yutalib.py

  Python library for data loading, saving and plotting.
  I think it's well commented.


pyezcalib.py
  in ./scripts/pylibs/pyezcalib.py

  Python library for ezca stuff.
  It has functions for recording and resetting default channel values in case of interrupt.


./scripts/PRCmodescan
  Python scripts for PRC modescan. Not well commented. Not organized.
  See elog #8012


./scripts/Alignment
  Python and shell scripts for alignment work. Not well commented.
  See elog #8164


./scripts/SUS/OplevCalibration
  Python scripts for oplev calibration. Not well commented.
  See elog #8221


./scripts/dither/gfactormeasurement
  Python scripts for g-factor measurement. Not well commented.
  See elog #8230


./scripts/SUS/ActuatorCalib
  Python scripts for calibrating actuators. Not well commented.
  See elog #8242

ELOG V3.1.3-